Systems of Equations
A chemist has two solutions of hydrochloric acid in stock: a 50% solution and an 80% solution. How many of each should she mix to obtain 100 milliliters of a 68% solution?

Answer by checkley77(12844)

.80X+.50(100-X)=.68*100
.80X+50-.50X=68
.30X=68-50
.30X=18
X=18/.30
X=60 ML OF 80% SOLUTION IS USED.
100-60=40 ML. OF 50% SOLUTION IS USED.
PROOF:
.80*60+.5*40=.68*100
48+20=68
68=68